Please answer in full with explanations, There are 2 separate parts.

Charlies Furniture Store has been in business for several years. The firms owners have described the store as a high-priceComplete this question by entering your answers in the tabs below. Required A Required B Required C Required D Required E Now
Answer #1

formulas used

sales margin = income/sales

turnover = sales / assets

ROI = margin / turnover

ans c

sales= assets * turnover

=1200,000 * 0.2 = $240,000

income = margin * sales

=27%*240,000 = $64,800

ROI = 27% / 20% = 1.35

now margin is reduced to 20%

keeping the ROI same

new sales = income / 20%

=64800/20% =$324000

thus the increase in sales is (324,000-240,000) = $84000

ans for part c

effect on margin is 20% as mentioned in part b. (decreased from 27%)

margin = income / sales 64800/324,000

effect on turnover

sales/ assets

324,000/1200,000 = 0.27 (increased from 0.2)

effect on ROI

margin / turnover

20%/ 0.27 = 0.74 (reduced from 1.34)